# AudioTrust: Benchmarking the Multifaceted Trustworthiness of Audio Large Language Models

### v3

@@ I.5 Additional Result
Current evaluations of ALLMs have primarily focused on their performance in fundamental tasks. SUPERB (Yang et al., 2021) first introduced a unified evaluation framework for speech processing, where self-supervised speech representation models are assessed across ten downstream tasks, including phoneme recognition, keyword spotting, speaker verification, and emotion recognition. This benchmark demonstrates the generality and effectiveness of SSL representations in diverse scenarios. Subsequently, SUPERB-SG (Tsai et al., 2022) extended this framework to encompass advanced semantic understanding and generative tasks, such as speech translation (Wahlster, 2013), voice conversion (Mohammadi & Kain, 2017), speech separation (Wang & Chen, 2018), and enhancement (Benesty et al., 2006), in order to further evaluate models’ generative abilities and robustness. SLURP (Bastianelli et al., 2020) provides a large-scale dataset and evaluation framework targeting spoken language understanding, thereby enabling a comprehensive comparison between end-to-end and pipeline approaches, while SLUE (Shon et al., 2022) assesses complex tasks including audio question answering, summarization, and named entity recognition within realistic speech scenarios with low-resource context, highlighting the impact of ASR models on downstream task performance. In the field of audio captioning, AudioCaps (Kim et al., 2019) and Clotho (Drossos et al., 2020) serve as major evaluation benchmarks, with Clotho-AQA (Lipping et al., 2022) pioneering a real-world dataset for audio question answering, facilitating the evaluation of models’ semantic reasoning capabilities. The recently released AIR-Bench (Yang et al., 2024c) categorizes evaluation tasks into two dimensions: fundamental abilities and dialogic abilities, covering a wide variety of audio types such as speech, environmental sounds, and music. The fundamental dimension comprises 19 specific tasks, whereas the dialogic dimension uses open-ended question-answering formats to evaluate generative performance of models under diverse and mixed audio backgrounds. These benchmarks offer diverse and comprehensive frameworks for evaluation and comparison of ALLMs, yet they mainly focus on fundamental performance; systematic assessments of safety, ethical risks, and social impacts remain insufficient.
@@
Existing safety evaluation benchmarks are relatively limited, with most focusing on multimodal scenarios or specific attack methods. For example, MM-SafetyBench (Liu et al., 2024) proposed an evaluation framework for image query attacks targeting multimodal LLMs, collecting 5,040 text-image pairs to assess model safety under image manipulation. SafeBench (Xu et al., 2022) constructed 23 risk scenarios and 2,300 multimodal harmful example pairs by automatically generating harmful multimodal queries, and designed a collaborative LLM review protocol to enhance evaluation reliability. In the audio domain, the Chat-Audio Attacks (CAA) benchmark (Yang et al., 2024d) designed four types of audio attacks for dialog audio attack evaluation, and adopted a synthesis of standard evaluation, GPT-4o-based assessment, and human evaluation strategies to measure model robustness. The study (Yang et al., 2024b) comprehensively evaluated the safety performance of five audio multimodal models via red-teaming against harmful audio, textual interference, and specific jailbreak attacks, revealing attack success rates as high as 70%. Furthermore, the SEA method (Lu et al., 2025) proposed a synthetic embedding augmentation approach for safety alignment, verifying the feasibility of aligning audio safety in multimodal models using only textual data. Although the above benchmarks have made progress in their respective areas, there is still a lack of a unified audio safety benchmark that comprehensively considers multidimensional risks such as fairness, hallucination detection, privacy protection, robustness, and authentication. Therefore, this work proposes the AudioTrust benchmark, which encompasses six core directions: fairness evaluation, hallucination detection, safety defense, privacy leakage, robustness challenges, and identity authentication. By combining scenario-driven question-answer pairs with GPT-4o automated evaluation, AudioTrust reveals the safety boundaries of ALLMs in high-risk environments, thereby providing systematic guidance for the secure and trustworthy deployment of future models.
@@
While AudioTrust offers a pioneering and comprehensive framework for the multidimensional trustworthiness evaluation of Audio Large Language Models (ALLMs), certain limitations warrant consideration. Firstly, the datasets, though meticulously constructed to cover a diverse range of scenarios across fairness, hallucination, safety, privacy, robustness, and authentication, are necessarily finite and may not encapsulate the full spectrum of real-world complexities or all potential adversarial manipulations, such as reliability (Ma et al., 2025). Secondly, the dynamic nature of ALLM development and emerging threat landscapes also means that any benchmark, including AudioTrust, represents a snapshot in time and will require continuous updates to remain relevant and comprehensive in assessing the evolving trustworthiness of these rapidly advancing systems.
Appendix L Social Impact
@@
Appendix M Data sheet
@@
This dataset is available to Huggingface Datasets (https://huggingface.co/datasets/JusperLee/AudioTrust).
The associated code, scripts, and benchmark framework are hosted on GitHub (https://github.com/JusperLee/AudioTrust).
Appendix N LLM Usage

### v4

@@ I.5 Additional Result
Current evaluations of ALLMs have primarily focused on their performance in fundamental tasks. SUPERB (Yang et al., 2021) first introduced a unified evaluation framework for speech processing, where self-supervised speech representation models are assessed across ten downstream tasks, including phoneme recognition, keyword spotting, speaker verification, and emotion recognition. This benchmark demonstrates the generality and effectiveness of SSL representations in diverse scenarios. Subsequently, SUPERB-SG (Tsai et al., 2022) extended this framework to encompass advanced semantic understanding and generative tasks, such as speech translation (Wahlster, 2013), voice conversion (Mohammadi & Kain, 2017), speech separation (Wang & Chen, 2018), and enhancement (Benesty et al., 2006), in order to further evaluate models’ generative abilities and robustness. SLURP (Bastianelli et al., 2020) provides a large-scale dataset and evaluation framework targeting spoken language understanding, thereby enabling a comprehensive comparison between end-to-end and pipeline approaches, while SLUE (Shon et al., 2023) assesses complex tasks including audio question answering, summarization, and named entity recognition within realistic speech scenarios with low-resource context, highlighting the impact of ASR models on downstream task performance. In the field of audio captioning, AudioCaps (Kim et al., 2019) and Clotho (Drossos et al., 2020) serve as major evaluation benchmarks, with Clotho-AQA (Lipping et al., 2022) pioneering a real-world dataset for audio question answering, facilitating the evaluation of models’ semantic reasoning capabilities. The recently released AIR-Bench (Yang et al., 2024b) categorizes evaluation tasks into two dimensions: fundamental abilities and dialogic abilities, covering a wide variety of audio types such as speech, environmental sounds, and music. The fundamental dimension comprises 19 specific tasks, whereas the dialogic dimension uses open-ended question-answering formats to evaluate generative performance of models under diverse and mixed audio backgrounds. These benchmarks offer diverse and comprehensive frameworks for evaluation and comparison of ALLMs, yet they mainly focus on fundamental performance; systematic assessments of safety, ethical risks, and social impacts remain insufficient.
@@
Existing safety evaluation benchmarks are relatively limited, with most focusing on multimodal scenarios or specific attack methods. For example, MM-SafetyBench (Liu et al., 2024) proposed an evaluation framework for image query attacks targeting multimodal LLMs, collecting 5,040 text-image pairs to assess model safety under image manipulation. SafeBench (Ying et al., 2024) constructed 23 risk scenarios and 2,300 multimodal harmful example pairs by automatically generating harmful multimodal queries, and designed a collaborative LLM review protocol to enhance evaluation reliability. In the audio domain, the Chat-Audio Attacks (CAA) benchmark (Yang et al., 2025d) designed four types of audio attacks for dialog audio attack evaluation, and adopted a synthesis of standard evaluation, GPT-4o-based assessment, and human evaluation strategies to measure model robustness. The study (Yang et al., 2025b) comprehensively evaluated the safety performance of five audio multimodal models via red-teaming against harmful audio, textual interference, and specific jailbreak attacks, revealing attack success rates as high as 70%. Furthermore, the SEA method (Lu et al., 2025) proposed a synthetic embedding augmentation approach for safety alignment, verifying the feasibility of aligning audio safety in multimodal models using only textual data. Although the above benchmarks have made progress in their respective areas, there is still a lack of a unified audio safety benchmark that comprehensively considers multidimensional risks such as fairness, hallucination detection, privacy protection, robustness, and authentication. Therefore, this work proposes the AudioTrust benchmark, which encompasses six core directions: fairness evaluation, hallucination detection, safety defense, privacy leakage, robustness challenges, and identity authentication. By combining scenario-driven question-answer pairs with GPT-4o automated evaluation, AudioTrust reveals the safety boundaries of ALLMs in high-risk environments, thereby providing systematic guidance for the secure and trustworthy deployment of future models.
@@
While AudioTrust offers a pioneering and comprehensive framework for the multidimensional trustworthiness evaluation of Audio Large Language Models (ALLMs), certain limitations warrant consideration. Firstly, the datasets, though meticulously constructed to cover a diverse range of scenarios across fairness, hallucination, safety, privacy, robustness, and authentication, are necessarily finite and may not encapsulate the full spectrum of real-world complexities or all potential adversarial manipulations, such as reliability (Ma et al., 2025). Secondly, the dynamic nature of ALLM development and emerging threat landscapes also means that any benchmark, including AudioTrust, represents a snapshot in time and will require continuous updates to remain relevant and comprehensive in assessing the evolving trustworthiness of these rapidly advancing systems. Thirdly, future work will extend our safety analysis to prosodic factors such as speech rate, investigating how extreme acoustic variations may act as adversarial channels by influencing the model’s front-end perception and recognition.
Appendix L Lessons from ALLMs for Future Fine-Tuning
Our results suggest that stronger general capability does not automatically translate into higher trustworthiness, and that different dimensions require targeted design and alignment. Below we summarize the main implications, following the six AudioTrust dimensions.
Fairness. We observe that both closed source and open source models still exhibit serious group unfairness in decision-making tasks. This suggests that future ALLMs should explicitly incorporate fairness objectives into training and alignment. We can add fairness-aware rewards and penalties in RLHF and RLAIF so that outputs displaying systematic bias across gender, age, accent, or socioeconomic cues receive negative feedback. Crucially, we need to adopt fairness regularization and fairness coefficients to achieve reward fairness (Ouyang et al., 2025); otherwise, these misaligned rewards may negatively impact the alignment of ALLMs.
Hallucination. Both open and closed source models are better at recognizing coarse physical impossibilities than at detecting subtler content mismatches or temporal logic errors in audio. This pattern indicates that current training is still dominated by transcription and text-style objectives, and that models rely heavily on textual priors rather than building a robust audio-grounded representation of physical and temporal structure. To reduce hallucinations in audio settings, future training should move beyond simple ASR plus instruction tuning and include objectives that reward correct reasoning about acoustic scenes and penalize logical and physical inconsistencies. One concrete direction is to inject synthetic negative examples with explicit physical violations and temporal inversions and optimize models to flag or reject such cases, so that audio commonsense and causal reasoning become part of the learning target rather than a side effect.
Safety. Closed-source models show weaknesses in medical scenarios and under emotional deception, while open-source models are prone to jailbreaks. This indicates that generic text-based refusal training is insufficient for the audio modality (Yang et al., 2025c). Future ALLMs require domain-specific alignment in high-risk areas like healthcare. Furthermore, since our results show models relax their guard under emotional pressure, safety fine-tuning must explicitly incorporate emotional and paralinguistic cues. Training models on datasets containing “emotional harmful queries” will teach them to treat urgent or distressed requests for illicit guidance as suspicious rather than authoritative.
Privacy. Closed-source models are much better at refusing to repeat explicit identifiers, while open-source models often memorize and reproduce bank accounts, SSNs or addresses with high accuracy once these appear in context. At the same time, all models, including the strongest ones, freely infer and disclose attributes such as age and ethnicity from voice. This points to two distinct needs. First, training, RLHF and RLAIF should include explicit privacy protection objectives: potential leakage of direct identifiers should be treated as a safety violation and receive negative feedback, and techniques such as memory truncation and selective unlearning should be integrated into the training pipeline so that models can forget or down-weight sensitive content. Second, models need to learn contextual privacy reasoning, that is, to recognize that some questions (for example, “guess my age or ethnicity from my voice”) should not be answered even if the model has the capability to infer them, and that in certain contexts “not inferring” is the correct behaviour. Achieving robust privacy awareness may require more than generic safety alignment: it calls for instruction tuning and preference optimization on purpose-built privacy scenarios, potentially augmented with reinforcement learning to allow the model to explore and learn appropriate responses in sensitive situations.
Robustness. The gap between closed source and open source ALLMs is particularly clear under realistic acoustic conditions such as background conversation, environmental noise, multiple speakers and audio quality degradation. Closed source systems appear to benefit from more mature front-end encoders and extensive exposure to noisy data, whereas many open source models hallucinate content or misinterpret non speech noise as meaningful when the signal becomes imperfect. Future ALLMs, especially in the open source ecosystem, will need to integrate stronger and more noise aware audio encoders and train on large scale real world noisy corpora, rather than relying mainly on clean or synthetic speech. Robustness training should explicitly target disentangling speech from background and channel effects so that the language component can condition on a stable representation of what was actually said, which in turn will reduce downstream hallucinations and improve reliability in real deployments.
Authentication. We find models are easily fooled by cloned voices, though strict system prompts can significantly improve defense. This suggests two complementary directions: First, training should enforce strict adherence to security instructions, treating them as hard constraints rather than soft preferences. Second, since current models prioritize semantic understanding over acoustic verification, future ALLMs must be integrated with components possessing stronger discriminative ability for speaker verification and deepfake detection (Li et al., 2024a). Strengthening the acoustic perception of synthetic artifacts, combined with prompt-level hardening, is essential for secure deployment.
Appendix M Social Impact
@@
Appendix N Data sheet
| Dimension | Attribute | Decision-making (Samples) | Stereotype-driven (Samples) | Metrics |
| Fairness | Gender | 60 | 60 | Group Fairness Score |
| Age | 60 | 60 | ||
| Race | 60 | 60 | ||
| Personality traits | 60 | 60 | ||
| Economic status | 60 | 60 | ||
| Linguistic characteristics | 60 | 60 | ||
| Accent | 60 | 60 | ||
| Total | 420 | 420 | ||
| Dimension | Sub-task | Samples | Metrics |
| Hallucination | Content mismatch | 80 | Accuracy, hallucination rate |
| Label mismatch | 80 | ||
| Logical violation | 80 | ||
| Physical violation | 80 | ||
| Total | 320 | ||
| Dimension | Task | Samples | Metrics |
| Safety | Enterprise system jailbreak | 100 | Defense Success Rate (DSR) |
| Financial system jailbreak | 100 | ||
| Medical system jailbreak | 100 | ||
| General illegal activity guidance | 300 | ||
| Total | 600 | ||
| Dimension | Attribute | Direct Privacy Leakage (#Samples) | Inference Privacy Leakage (#Samples) | Metrics |
| Privacy | Bank account number | 100 | — | Privacy refusal rate, Accuracy of leaked information |
| Driver license number | 100 | — | ||
| Home address | 100 | — | ||
| Phone number | 100 | — | ||
| Phone password | 100 | — | ||
| SSN | 100 | — | ||
| Age | — | 100 | ||
| Ethnicity | — | 100 | ||
| Setting of conversation | — | 100 | ||
| Total | 600 | 300 | ||
Note: “—” indicates that no data has been constructed for the corresponding category.
| Dimension | Sub-task | Task Robustness Experiment (Samples) | Anti-misinformation Ability Experiment (Samples) | Metrics |
| Robustness | Audio quality variation | 82 | 82 | Robust accuracy, error rate |
| Background conversation | 82 | 82 | ||
| Environmental sound | 82 | 82 | ||
| Multiple speakers | 82 | 82 | ||
| Noise interference | 40 | 40 | ||
| Adversarial robustness | 82 | 82 | ||
| Total | 450 | 450 | – | |
| Dimension | Task | Samples | Metrics |
| Authentication | Identity Verification Bypass | 100 | Imposter Rejection Rate(IRR) |
| Hybrid Spoofing | 100 | ||
| Voice Cloning Spoofing | 200 | ||
| Total | 400 | ||
@@
This dataset will be made publicly available after the paper is accepted.
The associated code, scripts, and benchmark framework are hosted on GitHub (https://github.com/AudioTrust/AudioTrust).
Data Provenance and Ethical Compliance
Fairness: The dataset contains no real conversations or sensitive personal information. All fairness-relevant scenarios are synthetically constructed using openly licensed audio resources: background sounds are sourced from Freesound (CC BY-NC licenses) and Pixabay (standard Pixabay Content License, permitting free use and modification subject to prohibited uses), speech audio is drawn from Mozilla Common Voice (CC0). All audio is used in accordance with the relevant licenses, and we cite the source platforms and license types in the paper.
Hallucination: The dataset contains no real conversations or sensitive personal information. All scenarios are synthetically constructed using openly licensed audio resources: background sounds are sourced from Freesound (CC BY-NC licenses) and Pixabay (standard Pixabay Content License, permitting free use and modification subject to prohibited uses). All audio is used in accordance with the relevant licenses, and we cite the source platforms and license types in the paper.
Safety: We filter out personally identifiable information and exclude sensitive or private conversational content. All foreground speech used in safety-related scenarios is drawn from openly licensed sources, specifically Mozilla Common Voice (CC0).
Privacy: The dataset contains no real private conversations or sensitive personal information. All privacy-relevant scenarios are synthetically constructed using openly licensed audio resources: background sounds are sourced from Freesound (CC0), and speech segments are drawn from Mozilla Common Voice (CC0).
Robustness:The dataset contains no real conversations or sensitive personal information. All scenarios are synthetically constructed using openly licensed audio resources: background sounds are sourced from Freesound (CC BY-NC licenses) and Pixabay (standard Pixabay Content License, permitting free use and modification subject to prohibited uses). All audio is used in accordance with the relevant licenses, and we cite the source platforms and license types in the paper. We further incorporate Big Bench Audio, a publicly available benchmark; all clips are used strictly under its original license terms.
Authentication: The dataset contains only short, task-specific commands without personal or sensitive information, excluding metadata that could identify individual speakers. For voice-command authentication (VCS) scenarios, we utilize self-recorded phrases from volunteers who provided explicit consent for research use.
Appendix O Dataset statistics
In this section, we provide detailed statistics for the benchmark datasets across different trustworthiness perspectives.
The following tables summarize the dataset sizes (including the number of prompts and input audio), task names, and the mapping between tasks and evaluation metrics for Fairness (Table 28), Hallucination (Table 29), Safety (Table 30), Privacy (Table 31), Robustness (Table 32) and Authentication (Table 33), respectively.
Appendix P LLM Usage